Find the area of an equilateral triangle whose perimeter is 36'.
Answer by checkley75(3666)

The reason for dividing the triangle into three equal parts (36/3) is because there are three equal sides to an eqilateral triangle. Without that piece of information there would be infinite solutions to this problem & I would not have attempted to help you!!!!!!!
---------------------------------------------------------------------
PERIMETER=36 THE THE SIDES=36/3=12'
THUS THE FORMULA IS B*H/2. TO FIND THE HEIGHT WE CONSTRUCT A RIGHT TRIANGLE WITH ONE SIDE (HYPOTENUSE)=12 & THE BASE IS 6 INCHES.
NOW WE FIND THE HEIGHT 12^2=6^2+X^2
144=36+X^2
X^2=144-36
X^2=108
X=10.39 FOR THE HEIGHT.
AREA=(6*10.39)/2
AREA=31.17